# Scheduling problem of aircrafts on a same runway and dual runways

### v5

@@ I Introduction
According to the studied objective functions, existing works on this topic can be mainly categorized into four optimization problems: the minimization problem of the total delay time [1]-[13], the minimization problem of the total deviation of scheduled takeoff and landing times [14]-[21], the minimization problem of the total makespan for arrival and departure aircraft [22]-[26], and the minimization problem of overall operational costs [27]-[29]. In existing works, the problems of rescheduling aircraft sequences were usually modeled as a mixed-integer optimization problem, which is essentially a NP-hard problem. Though the mixed-integer optimization models have wide applications and can find the global optimization solutions, the high computational complexity and the high computing time cost make the corresponding algorithms hard to completely match with desired performances.
@@
In this paper, the scheduling problems of landing and takeoff aircraft on a same runway and dual runways are addressed. In most of existing works, aircraft were considered based on the ICAO (International Civil Aviation Organization) separations, which includes $3$ classes of aircraft and there are totally $3\times 3=9$ kinds of pairwise combinations of landing and takeoff aircraft. In this paper, our work can be used to the RECAT-EU (Recovery Assistance for Cohesion and the Territories of Europe) systems, which might include $6$ or more classes of aircraft and there might be totally no smaller than $6\times 6=36$ kinds of pairwise combinations of landing and takeoff aircraft. The scheduling problems studied in this paper are much more comprehensive and complicated than the existing works based on ICAO separations. In contrast to the approaches based on mixed-integer optimization models in existing works, our approach focuses on the minimum separation times between aircraft by introducing some necessary assumptions and new concepts including relevance, breakpoint aircraft, path and class-monotonically-decreasing sequence. Four scheduling problems are discussed including landing scheduling problem, takeoff scheduling problem, and mixed landing and takeoff scheduling problems on a same runway and on dual runways. Two real-time optimal algorithms are proposed to find the optimal aircraft sequences to minimize the given objective function. Numerical examples are presented to show the effectiveness of the algorithms. In particular, when $100$ aircraft are considered, by using the algorithms in this paper, the optimal solution can be obtained in less than $5$ seconds, while by using the CPLEX software to solve the mix-integer optimization model, the optimal solution cannot be obtained within $1$ hour.
@@
Notations. The operation $A-B$ represents the set that consists of the elements of $A$ which are not elements of $B$; the operation $\phi_{0}-\phi_{1}$ represents the sequence which is obtained through modifying the sequence $\phi_{0}$ by removing the aircraft belonging to the sequence $\phi_{1}$ and keeping the orders of the rest aircraft unchanged; the symbol /// represents the meaning of “or”.
@@ II Problem formulation
where $Tf_{k}=[f_{k}^{\min},f_{k}^{\max}]$ represents the set of allowable takeoff or landing times, i.e., the time window constraint, for the aircraft $Tcf_{k}$ for two constant $f_{k}^{\min}\leq f_{k}^{\max}$, $Y_{ij}$ represents the minimum separation time between an aircraft $Tcf_{i}$ and its trailing aircraft $Tcf_{j}$. For the sake of expression convenience, when no confusion arises, $F(\phi,Sr(\phi),Pr_{0})$, $Sr(\phi)$ and $S_{i}(\phi)$ can be abbreviated as $F(\phi)$, $Sr$ and $S_{i}$.
@@
(Relevance) Consider an aircraft sequence $\phi=\langle Tcf_{1},Tcf_{2},\cdots,Tcf_{n}\rangle$. Let $S_{ij}=S_{j}-S_{i}$ represent the takeoff or landing separation time between an aircraft $Tcf_{i}$ and its trailing aircraft $Tcf_{j}$ for $i<j$. If $S_{ij}=Y_{{i}j}$, $i<j$, it is said that aircraft $Tcf_{j}$ is relevant to the aircraft $Tcf_{i}$.
@@
III Landing scheduling on a single runway
Let $T_{ij}$ represent the minimum separation time between an aircraft of class $i$ and a trailing aircraft of class $j$ on a single runway without considering the influence of other aircraft. Let $T_{0}=\min_{i,j\in\mathcal{I}}\{T_{ij}\}$ denote the minimum value of all possible separation times between aircraft, where $T_{0}$ is usually taken as $1$ minute.
Based on the minimum landing separation time standards at Heathrow Airport [30] and the understanding of the physical landing process, we propose the following assumptions.
@@ III Landing scheduling on a single runway
(2) For $i=\rho_{1},\rho_{2}$, $T_{ii}=T_{0}+\delta$, where $T_{0}/8<\delta<T_{0}/6$ is a positive integer, $\rho_{1}=3$ and $\rho_{2}=5$.
@@
(Resident-point aircraft) Consider a landing/takeoff sequence $\phi=\langle Tcf_{1},Tcf_{2},\cdots,Tcf_{n}\rangle$. If $S_{1}>t_{0}$, $Tcf_{1}$ is called a resident-point aircraft of $\phi$, and $S_{1}-t_{0}$ is called the resident time of $Tcf_{1}$. If the aircraft $Tcf_{i}$ is not relevant to the aircraft $Tcf_{i-1}$, i.e., $S_{(i-1)i}-Y_{(i-1)i}>0$, for $i=2,3,\cdots,n$, it is said that $Tcf_{i}$ is a resident-point of $\phi$ and $S_{(i-1)i}-Y_{(i-1)i}$ is the resident time of $Tcf_{i}$. Consider a mixed landing and takeoff sequence $\phi=\langle Tcf_{1},Tcf_{2},\cdots,Tcf_{n}\rangle$. Let $\mu_{1}$ and $\mu_{2}$ be the largest integers smaller than $i$ such that aircraft $Tcf_{\mu_{1}}$ is a landing aircraft and aircraft $Tcf_{\mu_{2}}$ is a takeoff aircraft. If the aircraft $Tcf_{i}$ is not relevant to the aircraft $Tcf_{\mu_{1}}$ and $Tcf_{\mu_{2}}$, $i=3,4,\cdots,n$, it is said that $Tcf_{i}$ is a resident-point aircraft of $\phi$ and $\min\{S_{\mu_{1}i}(\phi)-Y_{\mu_{1}{i}},S_{\mu_{2}i}(\phi)-Y_{\mu_{2}{i}}\}$ is the resident time of $Tcf_{i}$.
@@
Note that an aircraft $Tcf_{i}$ in a given sequence $\phi$ might be a breakpoint and a resident point at the same time.
@@
(3) Suppose that the sequence of landing aircraft is fixed. The optimization problem (6) is solved if and only if the aircraft $Tcf_{i+1}$ is relevant to the aircraft $Tcf_{i}$, $i=1,2,\cdots,n-1$.
@@
Remark 6
Theorem 1(3) shows that when the landing times of the aircraft have no constraints, the occurrence of resident-point aircraft should be avoided to ensure the optimality of the sequence.
@@
For convenience of expressions, under the condition of Assumption 5, $F(\phi,Sr(\phi))$ can be abbreviated as $F(\phi)$.
@@
In Lemmas 5 and 6, we discuss some typical cases for breakpoint aircraft which yields local minimum points.
@@
$\begin{array}[]{lll}F(\Phi_{a})-F(\Phi_{b})\\ =T_{cl_{s_{1}}cl_{s_{1}+1}}-T_{h_{1}cl_{s_{1}+1}}-T_{cl_{s_{1}}h_{1}}\\ +T_{h_{1}cl_{i+1}}+T_{cl_{i-1}h_{1}}-T_{cl_{i-1}cl_{i+1}}\\ <-0.5T_{0}-T_{0}+1.5T_{0}<0,\end{array}$
@@
$\begin{array}[]{lll}F(\Phi_{a})-F(\Phi_{b})\\ =T_{cl_{s_{1}}cl_{s_{1}+1}}-T_{h_{1}cl_{s_{1}+1}}-T_{cl_{s_{1}}h_{1}}+T_{h_{1}cl_{i+1}}\\ +T_{cl_{i-1}h_{1}}-T_{cl_{i-1}cl_{i+1}}\\ =T_{cl_{s_{1}}cl_{s_{1}+1}}-T_{h_{1}cl_{s_{1}+1}}-T_{cl_{s_{1}}h_{1}}+T_{h_{1}cl_{i+1}}\\ =T_{cl_{s_{1}}cl_{s_{1}+1}}-T_{h_{1}cl_{s_{1}+1}}+0.5T_{0},\end{array}$
@@
$\begin{array}[]{lll}F(\Phi_{a})-F(\Phi_{b})\\ \geq f_{1}(\Phi_{a})+f_{2}(\Phi_{a})-f_{1}(\Phi_{b})-f_{2}(\Phi_{b})-2\delta\\ \geq T_{cl_{n_{1}}cl_{n_{1}+1}}-1.5T_{0}-2\delta.\end{array}$
@@
$\begin{array}[]{lll}F(\Phi_{a})-F(\Phi_{b})\\ \geq f_{1}(\Phi_{a})+f_{2}(\Phi_{a})-f_{1}(\Phi_{b})-f_{2}(\Phi_{b})-T_{0}-2\delta\\ \geq T_{cl_{n_{1}}cl_{n_{1}+1}}-T_{0}-2\delta.\end{array}$
@@
Theorem 3 shows that when the set of allowable landing times for each aircraft is $Tf_{k}=[t_{0},+\infty]$, if all aircraft form a class-monotonically-decreasing sequence with no breakpoints, the optimization problem (6) can be solved.
From Theorems 1 and 2, it can be seen that the optimality of the aircraft sequence is heavily related to the resident-point aircraft, breakpoint aircraft and imaginary-breakpoint aircraft. To solve optimization problem (6), we can focus on these special aircraft so as to study the local minimum points of the objective function $F(\cdot)$. Due to the high complexity of the aircraft sequence and the constraints on the landing times of the aircraft, in the following, we only introduce some class-sequence sets, which can be calculated out offline, to study some typical scenarios and more general scenarios can be studied in a similar way.
@@
IV Takeoff scheduling on a single runway
@@
(3) Suppose that the takeoff aircraft sequence $\phi$ is fixed. The optimization problem (6) is solved if and only if $Tcf_{i+1}$ is relevant to $Tcf_{i}$, $i=1,2,\cdots,n-1$.
@@
In the following, we first give a calculation method for the objective function $F(\cdot)$ when the order of some aircraft is changed as discussed for landing sequeneces.
@@ V Scheduling problem of landing and takeoff aircraft on a same runway
In this section, the scheduling problem of mixed takeoff and landing aircraft on a same way for aircraft is discussed. Let $D_{T}$ denote the minimum separation time between a landing aircraft and a leading takeoff aircraft. Let $T_{D}$ denote the minimum separation time between a takeoff aircraft and a leading landing aircraft. As defined previously, $Y_{ij}$ is used to unifiedly denote the minimum separation time between any given two aircraft $Tcf_{i}$ and $Tcf_{j}$. Specifically, when a landing aircraft $Tcf_{j}$ and a leading takeoff aircraft $Tcf_{i}$ are considered, $Y_{ij}=D_{T}$; when a takeoff aircraft $Tcf_{j}$ and a leading landing aircraft $Tcf_{i}$ are considered, $Y_{ij}=T_{D}$; when two consecutive landing aircraft $Tcf_{i}$ and $Tcf_{j}$ are considered, $Y_{ij}=T_{cl_{i}cl_{j}}$; and when two consecutive takeoff aircraft $Tcf_{i}$ and $Tcf_{j}$ are considered, $Y_{ij}=D_{cl_{i}cl_{j}}$.
@@
From the previous sections, when the leading and the trailing aircraft are both landing or takeoff aircraft, the minimum separation time between an aircraft $Tcf_{i}$ and its leading aircraft is only relevant to their own classes. In contrast, when the takeoff and landing aircraft are simultaneously considered, the minimum separation time between an aircraft $Tcf_{i}$ and its leading aircraft might be related to not only their own classes, but also the classes of the aircraft ahead of them.
@@
When the scheduling problem of mixed landing and takeoff aircraft on a same runway, each aircraft might be relevant to two aircraft: one takeoff aircraft and one landing aircraft.
Remark 20
Theorem 6(3) shows that aircraft might be relevant to its nearest landing and takeoff aircraft ahead and is not relevant to other aircraft.
@@
Consider an aircraft sequence $\phi=\langle Tcf_{1},Tcf_{2},\cdots,Tcf_{n}\rangle$. Suppose that the aircraft sequence $\phi$ is fixed.
@@
Proof: (1) Sufficiency is obvious and we only discuss the necessity of this theorem. Suppose that there is no path from the aircraft $Tcf_{n}$ to the aircraft $Tcf_{1}$, when the optimization problem (6) is solved. Construct an aircraft set $\Theta_{1}$ such that there is a path from the aircraft $Tcf_{n}$ to the aircraft $Tcf_{i}$ for any $Tcf_{i}\in\Theta_{1}$ and there is no path from the aircraft $Tcf_{n}$ to the aircraft $Tcf_{i}$ for any $Tcf_{i}\in\{Tcf_{1},Tcf_{2},\cdots,Tcf_{n}\}-\Theta_{1}$. Shift the takeoff or landing times of all the aircraft in $\Theta_{1}$ as a whole to an earlier time instant with their relative landing or takeoff times unchanged while keeping necessary minimum separation times with the aircraft in $\{Tcf_{1},Tcf_{2},\cdots,Tcf_{n}\}-\Theta_{1}$. Such a operation can be realized because the aircraft in $\Theta_{1}$ are not relevant to any aircraft in $\{Tcf_{1},Tcf_{2},\cdots,Tcf_{n}\}-\Theta_{1}$. After this operation, the value of the objective function $F(\phi,Sr(\phi))$ becomes smaller, which yields a contraction. Therefore, there is a path from the aircraft $Tcf_{n}$ to the aircraft $Tcf_{1}$.
@@
Remark 21
From Theorem 7, the optimal value of the objective function $F(\cdot)$ can be calculated along the path from the aircraft $Tcf_{n}$ to the aircraft $Tcf_{1}$.
@@
Suppose that the aircraft sequence $\phi=\langle Tcf_{1},Tcf_{2},\cdots,Tcf_{n}\rangle$ is fixed. The aircraft $Tcf_{2}$ is relevant to aircraft $Tcf_{1}$, and aircraft $Tcf_{i+2}$ is relevant to aircraft $Tcf_{i+1}$ or aircraft $Tcf_{i}$, $i=1,2,\cdots,n-2$.
@@
Suppose that $Tf_{j}=[t_{0},+\infty]$ for all $j$, and Assumptions 1-4, 6-9 and 11-14 hold.
When the separation time of two consecutive takeoff or landing aircraft is large, by adding other landing or takeoff aircraft, the total operation time of the aircraft can be decreased. In the following, we make discussions on this issue.
@@
From Theorem 8(2), when the separation time between two takeoff or landing aircraft is large, takeoff-landing or landing-takeoff transition can be used to decrease the value of the objective function $F(\cdot)$.
Theorem 9
@@
In the proof of Theorem 9, what is different from the landing sequence and the takeoff sequence is that there is a landing-takeoff transition or a takeoff-landing transition at aircraft $Tcf_{n_{1}}$ in $\Phi_{a}$. Due to this reason, the term $D_{cl_{n_{1}}1}-Y_{n_{1}n_{2}}$ is included in the relationship between $F(\Phi_{a})$ and $F(\Phi_{b})$.
Remark 24
Theorem 9 gives a rule to calculate the objective function $F(\cdot)$ for scheduling problem of mixed takeoff and landing aircraft on a same runway.

### v7

# Scheduling problems of aircraft on a same runway and dual runways

@@ I Introduction
According to the studied objective functions, existing works on this topic can be mainly categorized into four optimization problems: the minimization problem of the total delay time [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13], the minimization problem of the total deviation of scheduled takeoff and landing times [14, 15, 16, 17, 18, 19, 20, 21], the minimization problem of the total makespan for arrival and departure aircraft [22, 23, 24, 25, 26], and the minimization problem of overall operational costs [27, 28, 29]. In existing works, the problems of rescheduling aircraft sequences were usually modeled as a mixed-integer optimization problem, which is essentially a NP-hard problem. Though the mixed-integer optimization models have wide applications and can find the global optimization solutions, the high computational complexity and the high computing time cost make the corresponding algorithms hard to completely match with desired performances.
@@
In this paper, the scheduling problems of landing and takeoff aircraft on a same runway and dual runways are addressed. In most of existing works, aircraft were considered based on the ICAO separations, which includes $3$ classes of aircraft and there are totally $3\times 3=9$ kinds of pairwise combinations of landing and takeoff aircraft. In this paper, our work can be used to the RECAT-EU systems, which might include $6$ or more classes of aircraft and there might be totally no smaller than $6\times 6=36$ kinds of pairwise combinations of landing and takeoff aircraft. The scheduling problems studied in this paper are much more comprehensive and complicated than the existing works based on ICAO separations. In contrast to the approaches based on mixed-integer optimization models in existing works, our approach focuses on the minimum separation times between aircraft by introducing some necessary assumptions and new concepts including relevance, breakpoint aircraft, path and class-monotonically-decreasing sequence. Four scheduling problems are discussed including landing scheduling problem, takeoff scheduling problem, and mixed landing and takeoff scheduling problems on a same runway and on dual runways. Two real-time optimal algorithms are proposed to find the optimal aircraft sequences to minimize the given objective function. Numerical examples are presented to show the effectiveness of the algorithms. In particular, when $100$ aircraft are considered, by using the algorithms in this paper, the optimal solution can be obtained in less than $5$ seconds, while by using the CPLEX software to solve the mix-integer optimization model, the optimal solution cannot be obtained within $1$ hour.
@@
Notations. The operation $A-B$ represents the set that consists of the elements of $A$ which are not elements of $B$; the operation $\phi_{0}-\phi_{1}$ represents the sequence which is obtained through modifying the sequence $\phi_{0}$ by removing the aircraft belonging to the sequence $\phi_{1}$ and keeping the orders of the rest aircraft unchanged; and the symbol /// represents the meaning of “or”.
@@ II Problem formulation
where $Tf_{k}=[f_{k}^{\min},f_{k}^{\max}]$ represents the set of allowable takeoff or landing times, i.e., the time window constraint, for the aircraft $Tcf_{k}$ for two constant $f_{k}^{\min}\leq f_{k}^{\max}$, $Y_{ij}$ represents the minimum separation time between an aircraft $Tcf_{i}$ and its trailing aircraft $Tcf_{j}$. For the sake of expression convenience, when no confusion arises, $F(\phi,Sr(\phi))$, $Sr(\phi)$ and $S_{i}(\phi)$ can be abbreviated as $F(\phi)$, $Sr$ and $S_{i}$.
@@
(Relevance) Consider an aircraft sequence $\phi=\langle Tcf_{1},Tcf_{2},\cdots,Tcf_{n}\rangle$. Let $S_{ij}=S_{j}-S_{i}$ represent the operation time interval between aircraft $Tcf_{i}$ and aircraft $Tcf_{j}$ for $i<j$. If $S_{ij}=Y_{{i}j}$ for $i<j$, it is said that aircraft $Tcf_{j}$ is relevant to the aircraft $Tcf_{i}$.
@@
III Landing scheduling problem on a single runway
Let $T_{ij}$ represent the minimum separation time between a landing aircraft of class $i$ and a trailing landing aircraft of class $j$ on a single runway without considering the influence of other aircraft. Let $T_{0}$ denote the minimum value of all possible separation times between aircraft, where $T_{0}$ is usually taken as $1$ minute.
Based on the minimum landing separation time standards at Heathrow Airport [30] and the understanding of the physical landing process, we propose the following assumptions and definitions.
@@ III Landing scheduling problem on a single runway
(2) For $i=\rho_{1},\rho_{2}$, $T_{ii}=T_{0}+\delta$, where $T_{0}/8<\delta<T_{0}/6$ is a positive number, $\rho_{1}=3$ and $\rho_{2}=5$.
@@
(Resident-point aircraft) Consider a landing/takeoff sequence $\phi=\langle Tcf_{1},Tcf_{2},\cdots,Tcf_{n}\rangle$. If $S_{1}>t_{0}$, $Tcf_{1}$ is called a resident-point aircraft of $\phi$, and $S_{1}-t_{0}$ is called the resident time of $Tcf_{1}$. If the aircraft $Tcf_{i}$ is not relevant to the aircraft $Tcf_{i-1}$, i.e., $S_{(i-1)i}-Y_{(i-1)i}>0$, for $i=2,3,\cdots,n$, it is said that $Tcf_{i}$ is a resident-point of $\phi$ and $S_{(i-1)i}-Y_{(i-1)i}$ is the resident time of $Tcf_{i}$. Consider a mixed landing and takeoff sequence $\phi=\langle Tcf_{1},Tcf_{2},\cdots,Tcf_{n}\rangle$. Let $\mu_{1}$ and $\mu_{2}$ be the largest integers smaller than $i$ for a landing aircraft and a takeoff aircraft $Tcf_{\mu_{2}}$. If the aircraft $Tcf_{i}$ is not relevant to the aircraft $Tcf_{\mu_{1}}$ and $Tcf_{\mu_{2}}$, $i=3,4,\cdots,n$, it is said that $Tcf_{i}$ is a resident-point aircraft of $\phi$ and $\min\{S_{\mu_{1}i}(\phi)-Y_{\mu_{1}{i}},S_{\mu_{2}i}(\phi)-Y_{\mu_{2}{i}}\}$ is the resident time of $Tcf_{i}$.
@@
Breakpoint aircraft and resident-point aircraft often exist in the actual aircraft sequence. It should be noted that an aircraft $Tcf_{i}$ in a given sequence $\phi$ might be a breakpoint and a resident point at the same time.
@@
In Theorem 1, we study the aircraft relevance and show that when the landing times of the aircraft have no constraints, the occurrence of resident-point aircraft should be avoided to ensure the optimality of the sequence.
@@
(3) Suppose that the orders of all landing aircraft in $\phi$ are fixed. The optimization problem (6) is solved if and only if the aircraft $Tcf_{i+1}$ is relevant to the aircraft $Tcf_{i}$, $i=1,2,\cdots,n-1$.
@@
In Lemmas 5 and 6, we discuss some typical cases for breakpoint aircraft which yields local minimum points. In Theorem 2, we give a rule to calculate the objective function $F(\cdot)$ based on a standard class-monotonically-decreasing sequence. In Theorem 3, we gives a condition for the optimality of the aircraft sequence.
Definition 4
@@
$\begin{array}[]{lll}&F(\Phi_{a})-F(\Phi_{b})\\ &=T_{cl_{s_{1}}cl_{s_{1}+1}}-T_{h_{1}cl_{s_{1}+1}}-T_{cl_{s_{1}}h_{1}}\\ &\quad+T_{h_{1}cl_{i+1}}+T_{cl_{i-1}h_{1}}-T_{cl_{i-1}cl_{i+1}}\\ &<-0.5T_{0}-T_{0}+1.5T_{0}<0,\end{array}$
@@
$\begin{array}[]{lll}&F(\Phi_{a})-F(\Phi_{b})\\ &=T_{cl_{s_{1}}cl_{s_{1}+1}}-T_{h_{1}cl_{s_{1}+1}}-T_{cl_{s_{1}}h_{1}}+T_{h_{1}cl_{i+1}}\\ &\quad+T_{cl_{i-1}h_{1}}-T_{cl_{i-1}cl_{i+1}}\\ &=T_{cl_{s_{1}}cl_{s_{1}+1}}-T_{h_{1}cl_{s_{1}+1}}-T_{cl_{s_{1}}h_{1}}+T_{h_{1}cl_{i+1}}\\ &=T_{cl_{s_{1}}cl_{s_{1}+1}}-T_{h_{1}cl_{s_{1}+1}}+0.5T_{0},\end{array}$
@@
$\begin{array}[]{lll}&F(\Phi_{a})-F(\Phi_{b})\\ &\geq f_{1}(\Phi_{a})+f_{2}(\Phi_{a})-f_{1}(\Phi_{b})-f_{2}(\Phi_{b})-2\delta\\ &\geq T_{cl_{n_{1}}cl_{n_{1}+1}}-1.5T_{0}-2\delta.\end{array}$
@@
$\begin{array}[]{lll}&F(\Phi_{a})-F(\Phi_{b})\\ &\geq f_{1}(\Phi_{a})\!+f_{2}(\Phi_{a})\!-f_{1}(\Phi_{b})\!-f_{2}(\Phi_{b})-T_{0}-2\delta\\ &\geq T_{cl_{n_{1}}cl_{n_{1}+1}}-T_{0}-2\delta.\end{array}$
@@
Theorem 3 shows that when the set of allowable landing times for each aircraft is $Tf_{k}=[t_{0},+\infty]$, if all aircraft form a class-monotonically-decreasing sequence without breakpoints, the optimization problem (6) can be solved.
From Theorems 1 and 2, it can be seen that the optimality of the aircraft sequence is heavily related to the resident-point aircraft, breakpoint aircraft and the aircraft with separation times larger than $T_{0}$. To solve optimization problem (6), we can focus on these special aircraft so as to study the local minimum points of the objective function $F(\cdot)$. Due to the high complexity of the aircraft sequence and the constraints on the landing times of the aircraft, in the following, we only introduce some class-sequence sets, which can be calculated out offline, to study some typical scenarios and more general scenarios can be studied in a similar way.
@@
IV Takeoff scheduling problem on a single runway
@@
As discussed in Theorem 1, in Theorem 4, we study the aircraft relevance and show that when the takeoff times of the aircraft have no constraints, the occurrence of resident-point aircraft should be avoided to ensure the optimality of the sequence.
@@
(3) Suppose that the orders of all the takeoff aircraft in $\phi$ is fixed. The optimization problem (6) is solved if and only if $Tcf_{i+1}$ is relevant to $Tcf_{i}$, $i=1,2,\cdots,n-1$.
@@
In the following, we first give a calculation method for the objective function $F(\cdot)$ when the order of some aircraft is changed as discussed for landing sequences.
@@
Theorem 6
Consider a group of takeoff aircraft $\{Tcf_{1},Tcf_{2},\cdots,Tcf_{n}\}$, where $cl_{1}\geq cl_{2}\geq cl_{3}\geq\cdots\geq cl_{n}$. Suppose that Assumptions 6-10 hold for each takeoff aircraft sequence. Then the optimization problem (6) can be solved if the takeoff aircraft sequence is taken as $\phi_{0}=\langle Tcf_{1},Tcf_{2},\cdots,Tcf_{n}\rangle$.
@@ V Scheduling problem of landing and takeoff aircraft on a same runway
In this section, the scheduling problem of mixed takeoff and landing aircraft on a same way is discussed. Let $D_{T}$ denote the minimum separation time between a landing aircraft and a leading takeoff aircraft. Let $T_{D}$ denote the minimum separation time between a takeoff aircraft and a leading landing aircraft. As defined previously, $Y_{ij}$ is used to unifiedly denote the minimum separation time between any given two aircraft $Tcf_{i}$ and $Tcf_{j}$. Specifically, when a landing aircraft $Tcf_{j}$ and a leading takeoff aircraft $Tcf_{i}$ are considered, $Y_{ij}=D_{T}$; when a takeoff aircraft $Tcf_{j}$ and a leading landing aircraft $Tcf_{i}$ are considered, $Y_{ij}=T_{D}$; when two consecutive landing aircraft $Tcf_{i}$ and $Tcf_{j}$ are considered, $Y_{ij}=T_{cl_{i}cl_{j}}$; and when two consecutive takeoff aircraft $Tcf_{i}$ and $Tcf_{j}$ are considered, $Y_{ij}=D_{cl_{i}cl_{j}}$.
@@
From the previous sections, when the leading and the trailing aircraft are both landing or takeoff aircraft, the minimum separation time between an aircraft and its leading aircraft is only relevant to their own classes. In contrast, when the takeoff and landing aircraft are simultaneously considered, the minimum separation time between an aircraft and its leading aircraft might be related to not only their own classes, but also the classes of the aircraft ahead of them.
In Theorem 7, we study the aircraft relevance and show that when the operation times of the aircraft have no constraints, each trailing aircraft should be relevant to at least one of its nearest landing and takeoff aircraft ahead to ensure the optimality of the sequence. Theorem 8 shows that the optimal value of the objective function $F(\cdot)$ can be calculated along the path from the last aircraft to the first aircraft.
@@
When the scheduling problem of mixed landing and takeoff aircraft on a same runway, each aircraft might be relevant to two aircraft ahead: one takeoff aircraft and one landing aircraft, and is not relevant to other aircraft.
@@
Consider an aircraft sequence $\phi=\langle Tcf_{1},Tcf_{2},\cdots,Tcf_{n}\rangle$. Suppose that the orders of all the aircraft in $\phi$ is fixed.
@@
Proof: (1) Sufficiency is obvious and we only discuss the necessity of this theorem. Suppose that there is no path from the aircraft $Tcf_{n}$ to the aircraft $Tcf_{1}$, when the optimization problem (6) is solved. Construct an aircraft set $\Theta_{1}$ such that there is a path from the aircraft $Tcf_{n}$ to the aircraft $Tcf_{i}$ for any $Tcf_{i}\in\Theta_{1}$ and there is no path from the aircraft $Tcf_{n}$ to the aircraft $Tcf_{i}$ for any $Tcf_{i}\in\{Tcf_{1},Tcf_{2},\cdots,Tcf_{n}\}-\Theta_{1}$. Clearly, $Tcf_{1}\notin\Theta_{1}$. Shift the takeoff or landing times of all the aircraft in $\Theta_{1}$ as a whole to an earlier time instant with their relative landing or takeoff times unchanged while keeping necessary minimum separation times with the aircraft in $\{Tcf_{1},Tcf_{2},\cdots,Tcf_{n}\}-\Theta_{1}$. Such an operation can be realized because the aircraft in $\Theta_{1}$ are not relevant to any aircraft in $\{Tcf_{1},Tcf_{2},\cdots,Tcf_{n}\}-\Theta_{1}$. After this operation, the value of the objective function $F(\phi,Sr(\phi))$ becomes smaller, which yields a contraction. Therefore, there is a path from the aircraft $Tcf_{n}$ to the aircraft $Tcf_{1}$.
@@
Consider the aircraft sequence $\phi=\langle Tcf_{1},Tcf_{2},\cdots,Tcf_{n}\rangle$. Suppose that the aircraft $Tcf_{2}$ is relevant to aircraft $Tcf_{1}$, and aircraft $Tcf_{i+2}$ is relevant to aircraft $Tcf_{i+1}$ or aircraft $Tcf_{i}$, $i=1,2,\cdots,n-2$.
@@
Suppose that $Tf_{j}=[t_{0},+\infty]$ for all $j$, and Assumptions 1-4, 6-9 and 11-13 hold.
When the separation time of two consecutive takeoff/landing aircraft is large, by adding one aircraft with different operation task, the total operation time of the aircraft can be decreased. In the following, we make discussions about this issue.
@@
In Theorem 10, what is different from the landing sequence and the takeoff sequence is that there is a landing-takeoff transition or a takeoff-landing transition at aircraft $Tcf_{n_{1}}$ in $\Phi_{a}$. Due to this reason, the term $D_{cl_{n_{1}}1}-Y_{n_{1}n_{2}}$ is included in the relationship between $F(\Phi_{a})$ and $F(\Phi_{b})$.
@@
Theorem 10 gives a rule to calculate the value of the objective function $F(\cdot)$ for scheduling problem of mixed takeoff and landing aircraft on a same runway.